# Why We Need More than General Multimodal LLMs

### v1

@@ Abstract
While recent Multimodal Large Language Models exhibit impressive capabilities for general multimodal tasks, specialized domains like music necessitate tailored approaches. Music Audio-Visual Question Answering (Music AVQA) particularly underscores this, presenting unique challenges with its continuous, densely layered audio-visual content, intricate temporal dynamics, and the critical need for domain-specific knowledge. Through a systematic analysis of Music AVQA datasets and methods, this position paper identifies that specialized input processing, architectures incorporating dedicated spatial-temporal designs, and music-specific modeling strategies are critical for success in this domain. Our study provides valuable insights for researchers by highlighting effective design patterns empirically linked to strong performance, proposing concrete future directions for incorporating musical priors, and aiming to establish a robust foundation for advancing multimodal musical understanding. This work is intended to inspire broader attention and further research, supported by a continuously updated anonymous GitHub repository of relevant papers: https://github.com/xid32/Survey4MusicAVQA.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive efficacy across a wide range of tasks, modelling various modalities such as text, image, and audio girdhar2023imagebind ; han2024onellm ; bai2025qwen2 ; hurst2024gpt ; liang2024querying ; huang2024audiogpt ; diao2025temporal . However, this success brings forth a critical consideration: the tension between the broad applicability of general multimodal approaches and the requirements of specialized domains 8665366 ; weck2024muchomusic ; liu2024music . This leads to a central question: Are general-purpose MLLMs, despite their advancements, truly sufficient for all multimodal tasks, especially those demanding deep, domain-specific understanding?
Music Performance Audio-Visual Question Answering (Music AVQA) emerges as a particularly challenging multimodal domain that compellingly illustrates this tension 9879157 ; 10484352 ; jiang-yin-2023-target ; ye2024cat . Music, in its rich complexity, often requires specialized treatment beyond the capabilities of generic models diao2024learning ; le2025natural ; bogdanov2019mtg ; chu2023qwen ; deng2023musilingo ; yuan2024chatmusician ; zhao2024openmu ; agostinelli2023musiclm ; lu2023musecoco . Unlike common scenarios with sparse and discrete audio signals, music performances exhibit a continuous and tightly interwoven blend of audio and visual signals, offering a uniquely rich context for fine-grained audio-visual scene understanding and temporal reasoning 9879157 ; 10484352 ; diao2024learning ; ma2024look ; li2024object ; li2023progressive . For instance, tasks such as discerning the loudest instrument amidst an ensemble or comparing rhythmic complexity between two spatially distinct performers demand a level of granularity that general-purpose MLLMs may not inherently possess liu2024music ; diao2024learning ; li2024object . Thus, Music AVQA serves as an ideal lens through which to examine the limitations of general multimodal approaches and to advocate for the necessity of domain-specific adaptations 9879157 ; diao2024learning .
@@ 2 Background
Solo Performance – A single musician showcasing technical skills and artistic expression on one instrument. 2 Ensemble of the Same Instrument – Multiple musicians playing identical or related instruments, creating unified harmonies and textures. 3 Ensemble of Different Instruments – Musicians performing with a variety of instruments, producing diverse tonal colors and complex musical interactions. 4 Culture-Specific Ensemble – Traditional instrumental groups that embody the musical heritage and regional styles of specific cultures. See Appendix Section B for examples.
@@
Existential Questions: Determine whether a sound corresponds to a visible object in the scene (e.g., “Is this sound from the instrument in the video?"). 2 Counting Questions: Quantify audio-visual elements that require cross-modal integration (e.g., “How many instruments are sounding in the video?"). 3 Location Questions: Identify the spatial position of sound sources within the visual scene (e.g., “Where is the first sounding instrument?"). 4 Comparative Questions: Compare properties across different audio-visual elements (e.g., “Is the instrument on the left louder than the one on the right?"). 5 Temporal Questions: Reason about the timing and sequential relationships between auditory and visual events (e.g., “Which instrument produces sound before the piano?"). See Appendix Section C for examples.
@@ 3 Evolution of MUSIC-AVQA Datasets
The development of Music AVQA research has been driven by progressively refined datasets addressing specific limitations. As summarized in Table 6 in Appendix Section E, this evolution began with the 1 MUSIC-AVQA dataset 9879157 , the first large-scale benchmark designed specifically for AVQA in musical contexts, comprising 9,288 performance videos and 45,867 question-answer pairs across diverse reasoning tasks. Subsequent research reveal challenges related to data bias and imbalanced answer distributions, prompting the creation of 2 MUSIC-AVQA v2.0 10484352 , which expands to 10,518 videos and approximately 54,000 question-answer pairs. This version balance 15 biased templates by ensuring no dominant answers exceed 60% for binary questions or 50% for multi-class questions, particularly enhancing representation in various question categories. Building on these foundations, 3 MUSIC-AVQA-R ma2024look introduces robustness evaluation through question rephrasing, expanding the test set from 9,129 to 211,572 questions. With a vocabulary five times larger than the original dataset, MUSIC-AVQA-R distinguishes between head (common) and tail (rare) samples, enabling assessment of model performance in both in-distribution and out-of-distribution scenarios. This progressive refinement of datasets has laid a solid foundation for advancing multimodal understanding and robust evaluation in music performance environments.
Table 1: Architectural summary of representative Music AVQA methods. Each method lists the text, visual, and audio encoders used, along with an indication of whether explicit spatial-temporal (S-T) modeling is incorporated. Detailed descriptions of each method are provided in Appendix F and G.
@@
Amuse diao2024learning Transformer NIPS2017_3f5ee243 Swin-Transformer-v2 liu2022swin HTS-AT chen2022hts
Audio Flamingo 10.5555/3692070.3693076 OPT-IML-MAX-1.3B iyer2023optimlscalinglanguagemodel - ClapCap 10448504
AVMoE cheng2025mixtures - Swin-Transformer-v2 liu2022swin HTS-AT chen2022hts $\times$
AVSD schwartz2019simple LSTM LSTM LSTM $\times$
AVSiam lin2024siamese - ViT radford2021learning ViT radford2021learning $\times$
AVST 9879157 LSTM ResNet-18 he2016deep VGGish 45857
CAT ye2024cat LLaMA2-7B touvron2023llama ImageBind girdhar2023imagebind ImageBind girdhar2023imagebind $\times$
ChatBridge zhao2023chatbridge Vicuna-13B vicuna2023 ViT-G sun2023evaclipimprovedtrainingtechniques BEATs $1$$\times$
CIGN mo2023class - ResNet-18 he2016deep ResNet-18 he2016deep
COCA lao2023coca Word Embedding ResNet-18 he2016deep VGGish 45857 $\times$
CONVLSTM 9145807 LSTM - Conv $\times$
CrossMAE guo2024crossmae - MAE 9879206 AudioMAE 10.5555/3600270.3602351 $\times$
DCL lv2023disentangled DeBERTa-V3-Large he2021debertav3 ViT radford2021learning AST gong2021ast
DG-SCT duan2023cross - ViT radford2021learning HTS-AT chen2022hts
EEMC 10.1007/978-3-031-72904-1_12 RoBERTa zhuang-etal-2021-robustly ViT radford2021learning VGGish 45857
FCNLSTM 9145807 LSTM - Conv $\times$
GPT-4o hurst2024gpt Transformer CLIP-ViT Transformer $\times$
GRU 7410636 LSTM VGGNet simonyan2014very - $\times$
HCRN le2020hierarchical BiLSTM ResNet-18 he2016deep - $\times$
LAST-Att 10484352 LSTM Swin-Transformer-v2 liu2022swin Audio-Spectrogram-Transformer
LAVisH lin2023vision - ViT radford2021learning ViT radford2021learning
LAViT yun2021pano Transformer NIPS2017_3f5ee243 Transformer NIPS2017_3f5ee243 Transformer NIPS2017_3f5ee243
LSTTA liu2023parameter CLIP radford2021learning CLIP radford2021learning w2v-Conformer gulati2020conformer
MAVEN ma2025fortisavqa Mixtral InternViT-300M-448px chen2024internvl Transformer $\times$
MCAN 8953581 GloVe pennington-etal-2014-glove +LSTM Faster R-CNN NIPS2015_14bfa6bb - $\times$
MCCD ma2024look - - -
Meerkat 10.1007/978-3-031-73039-9_4 LLaMA2-7B touvron2023llama CLIP-ViT CLAP 10095889
OGM wei2024fly - ResNet-18 he2016deep ResNet-18 he2016deep $\times$
@@
OneLLM han2024onellm LLaMA2-7B touvron2023llama CLIP-ViT Unified Multimodal Encoder $\times$
OPM wei2024fly - ResNet-18 he2016deep ResNet-18 he2016deep $\times$
PSAC 10.1609/aaai.v33i01.33018658 Word Embedding CNN - $\times$
PSTP-Net li2023progressive CLIP radford2021learning CLIP radford2021learning VGGish 45857
@@
QaP liang2024querying DeBERTa-V2-XLarge CLIP radford2021learning CLAP 10095889 $\times$
Qwen2.5-VL bai2025qwen2 MRoPE bai2025qwen2 ViT radford2021learning - $\times$
RefAtomNet peng2024referring BERT ViT radford2021learning -
VALOR 10.1109/TPAMI.2024.3479776 BERT CLIP radford2021learning AST gong2021ast $\times$
VAST 10.5555/3666122.3669307 BERT devlin2019bert ViT dosovitskiy2020image BEATs $1$$\times$
VideoLLaMA-2 cheng2024videollama Transformer CLIP radford2021learning BEATs 10.5555/3618408.3618611
VITA fu2024vita Mixtral jiang2024mixtral InternViT-300M-448px chen2024internvl CNN $\times$
@@ 4 Categorization of Music AVQA Methods Based on Architecture
CNN-based models. CNN-based methods typically utilize convolutional backbones such as ResNet or VGGish to encode modality-specific information into global or regional features, often relying on simpler late-stage fusion strategies. The AVST method exemplifies this approach, combining ResNet-18 visual embeddings and VGGish audio features through spatial attention modules to explicitly localize sound sources within visual frames. PSTP-Net extends this design by introducing a progressive refinement strategy that sequentially filters temporal segments and spatial regions, systematically narrowing down question-relevant audio-visual content prior to fusion. Although CNN-based models are computationally efficient and straightforward, their reliance on late fusion may pose challenges to capturing the complex temporal dynamics characteristic of musical performances.
Hybrid models. Hybrid models combine CNNs, transformers, and large language models (LLMs) to enable unified multimodal reasoning. They typically employ pre-trained encoders from both CNN and transformer families, integrated through sophisticated cross-modal fusion mechanisms. Representative examples include ChatBridge, CAT, OneLLM, and Meerkat. ChatBridge utilizes a perceiver-based multimodal transformer to merge modalities via language-aligned latent representations, followed by a frozen LLM for reasoning. CAT introduces modality-specific clue aggregation modules on top of ImageBind encodings, enabling precise question-driven multimodal grounding before passing information to a generative LLaMA2 LLM. OneLLM further generalizes multimodal integration by introducing a universal projection mechanism that allows a single LLM to interpret diverse modality embeddings seamlessly. In contrast, Meerkat emphasizes fine-grained cross-modal alignment through an audio-visual optimal transport module that explicitly matches audio segments to corresponding visual regions, achieving strong performance on tasks requiring precise localization of sound sources, underscoring the benefit of precise local grounding for complex audio-visual interactions in musical contexts.
@@ 5 A Call on Specialized Multimodal Input Processing for Music AVQA
Compared to event-centric AVQA tasks that typically involve short, discrete sound events and lower-resolution recordings, Music AVQA deals with continuous, polyphonic streams spanning multiple spatial and temporal scales. Audio is commonly sampled at high rates (44.1 kHz or above) and often preserved in lossless formats to retain subtle timbral and articulatory detail slizovskaia2021conditioned . Visual inputs similarly tend to require higher resolution (1080p or above) and frame rates (30–60 fps) to capture nuanced performer motions such as bowing or fingering gan2020music ; jin2024audio . Even modest temporal offsets—around 100–200ms—can affect the perceived correspondence between gesture and sound. To improve synchronization and cue isolation, some recent models adopt preprocessing strategies like beat-based segmentation 10.1145/2964284.2973795 and harmonic-percussive separation fitzgerald2010harmonic , which can help surface rhythmically or acoustically meaningful content for downstream reasoning.
@@
Tokenization plays a central role in structuring inputs for multimodal reasoning, and recent Music AVQA models often tailor their strategies to preserve musical structure. For audio, models such as Amuse diao2024learning , DG-SCT duan2023cross , and PSTP-Net li2023progressive transform waveforms into Mel-spectrograms, which are then segmented via patch-based encoders like AST gong2021ast and HTS-AT chen2022hts or CNNs such as VGGish 45857 and ResNet-18 he2016deep . Audio Flamingo 10.5555/3692070.3693076 , for instance, uses overlapping 7-second windows in ClapCap 10448504 to embed long-range audio context. Visual streams are frequently tokenized using ViT radford2021learning or Swin-based liu2022swin patch embeddings (e.g., in AVSiam lin2024siamese and LAVisH lin2023vision ), while earlier models like AVST 9879157 use frame-level CNN features. Text tokenization is typically handled by subword models aligned with large language models (e.g., LLaMA2 touvron2023llama , RoBERTa zhuang-etal-2021-robustly ), as seen in ChatBridge zhao2023chatbridge and OneLLM han2024onellm . These tokenization schemes help preserve temporal granularity and modality alignment, which may be important for interpreting overlapping instruments, rhythmic changes, and localized visual cues.
@@
Music AVQA tasks often involve long-form performances with overlapping sources and evolving musical dynamics, which can create challenges for segmentation, annotation, and generalization. Unlike typical AVQA datasets centered on short clips and isolated actions, music-focused benchmarks (e.g., MUSIC-AVQA 9879157 ) include multi-instrument performances spanning several minutes. These conditions place greater demands on dataset diversity to avoid overfitting to genre-specific patterns or ensemble configurations. To broaden coverage, some models are trained on data drawn from live performances, studio recordings, and synthetic renderings. However, the absence of symbolic structure can limit the model’s access to mid-level grounding. In this context, musically informed preprocessing (e.g., onset alignment, rhythmic segmentation, graph representation learning huang2024exploring ) may support more interpretable and temporally aligned input representations.

### v2

# Music Audio-Visual Question Answering Requires Specialized Multimodal Designs

@@ Abstract
While recent Multimodal Large Language Models exhibit impressive capabilities for general multimodal tasks, specialized domains like music necessitate tailored approaches. Music Audio-Visual Question Answering (Music AVQA) particularly underscores this, presenting unique challenges with its continuous, densely layered audio-visual content, intricate temporal dynamics, and the critical need for domain-specific knowledge. Through a systematic analysis of Music AVQA datasets and methods, this paper identifies that specialized input processing, architectures incorporating dedicated spatial-temporal designs, and music-specific modeling strategies are critical for success in this domain. Our study provides valuable insights for researchers by highlighting effective design patterns empirically linked to strong performance, proposing concrete future directions for incorporating musical priors, and aiming to establish a robust foundation for advancing multimodal musical understanding. We aim to encourage further research in this area and provide a GitHub repository of relevant works: https://github.com/WenhaoYou1/Survey4MusicAVQA.
Music Audio-Visual Question Answering Requires Specialized Multimodal Designs
Wenhao You∗1, Xingjian Diao∗2, Wenjun Huang3, Chunhui Zhang2, Keyi Kong2, Weiyi Wu2, Chiyu Ma2, Zhongyu Ouyang2, Tingxuan Wu4, Ming Cheng2, Soroush Vosoughi2, Jiang Gui2 1University of Waterloo, 2Dartmouth College, 3UC Irvine, 4New York University w22you@uwaterloo.ca, xingjian.diao.gr@dartmouth.edu
WY and XD contributed equally and are listed alphabetically by their first name. WY, from University of Waterloo, was mentored by XD from Dartmouth College.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive efficacy across a wide range of tasks, modelling various modalities such as text, image, and audio Girdhar et al. (2023); Han et al. (2024a); Bai et al. (2025); Hurst et al. (2024); Liang et al. (2024); Huang et al. (2024b); Diao et al. (2025d). However, this success brings forth a critical consideration: the tension between the broad applicability of general multimodal approaches and the requirements of specialized domains Simonetta et al. (2019); Weck et al. (2024); Liu et al. (2024a). This leads to a central question: Are general-purpose MLLMs, despite their advancements, truly sufficient for all multimodal tasks, especially those demanding deep, domain-specific understanding?
Music Performance Audio-Visual Question Answering (Music AVQA) emerges as a particularly challenging multimodal domain that compellingly illustrates this tension Li et al. (2022); Liu et al. (2024b); Jiang and Yin (2023); Ye et al. (2024). Music, in its rich complexity, often requires specialized treatment beyond the capabilities of generic models Le et al. (2025); Bogdanov et al. (2019); Chu et al. (2023); Diao et al. (2025c); Deng et al. (2023); Yuan et al. (2024); Zhao et al. (2024); Agostinelli et al. (2023); Lu et al. (2023). Unlike common scenarios with sparse and discrete audio signals, music performances exhibit a continuous and tightly interwoven blend of audio and visual signals, offering a uniquely rich context for fine-grained audio-visual scene understanding and temporal reasoning Li et al. (2022); Liu et al. (2024b); Diao et al. (2024); Ma et al. (2024); Li et al. (2024c, 2023). For instance, tasks such as discerning the loudest instrument amidst an ensemble or comparing rhythmic complexity between two spatially distinct performers demand a level of granularity that general-purpose MLLMs may not inherently possess Liu et al. (2024a); Diao et al. (2024); Li et al. (2024c). Thus, Music AVQA serves as an ideal lens through which to examine the limitations of general multimodal approaches and to advocate for the necessity of domain-specific adaptations Li et al. (2022); Diao et al. (2024).
@@ 2 Background
Solo Performance – A single musician showcasing technical skills and artistic expression on one instrument. 2 Ensemble of the Same Instrument – Multiple musicians playing identical or related instruments, creating unified harmonies and textures. 3 Ensemble of Different Instruments – Musicians performing with a variety of instruments, producing diverse tonal colors and complex musical interactions. 4 Culture-Specific Ensemble – Traditional instrumental groups that embody the musical heritage and regional styles of specific cultures. See Appendix Section B for detailed examples.
@@
Existential Questions: Determine whether a sound corresponds to a visible object in the scene (e.g., “Is this sound from the instrument in the video?"). 2 Counting Questions: Quantify audio-visual elements that require cross-modal integration (e.g., “How many instruments are sounding in the video?"). 3 Location Questions: Identify the spatial position of sound sources within the visual scene (e.g., “Where is the first sounding instrument?"). 4 Comparative Questions: Compare properties across different audio-visual elements (e.g., “Is the instrument on the left louder than the one on the right?"). 5 Temporal Questions: Reason about the timing and sequential relationships between auditory and visual events (e.g., “Which instrument produces sound before the piano?"). See Appendix Section C for detailed examples.
@@ 3 Evolution of MUSIC-AVQA Datasets
The development of Music AVQA research has been driven by progressively refined datasets addressing specific limitations. As summarized in Table 6 in Appendix Section E, this evolution began with the 1 MUSIC-AVQA dataset Li et al. (2022), the first large-scale benchmark designed specifically for AVQA in musical contexts, comprising 9,288 performance videos and 45,867 question-answer pairs across diverse reasoning tasks. Subsequent research reveal challenges related to data bias and imbalanced answer distributions, prompting the creation of 2 MUSIC-AVQA v2.0 Liu et al. (2024b), which expands to 10,518 videos and approximately 54,000 question-answer pairs. This version balance 15 biased templates by ensuring no dominant answers exceed 60% for binary questions or 50% for multi-class questions, particularly enhancing representation in various question categories. Building on these foundations, 3 MUSIC-AVQA-R Ma et al. (2024) introduces robustness evaluation through question rephrasing, expanding the test set from 9,129 to 211,572 questions. With a vocabulary five times larger than the original dataset, MUSIC-AVQA-R distinguishes between head (common) and tail (rare) samples, enabling assessment of model performance in both in-distribution and out-of-distribution scenarios. This progressive refinement of datasets has laid a solid foundation for advancing multimodal understanding and robust evaluation in music performance.
@@
Amuse Diao et al. (2024) Transformer Vaswani et al. (2017) Swin-Transformer-v2 Liu et al. (2022) HTS-AT Chen et al. (2022)
Audio Flamingo Kong et al. (2024) OPT-IML-MAX-1.3B Iyer et al. (2022) - ClapCap Elizalde et al. (2024)
AVMoE Cheng et al. (2025) - Swin-Transformer-v2 Liu et al. (2022) HTS-AT Chen et al. (2022) $\times$
AVSD Schwartz et al. (2019) LSTM LSTM LSTM $\times$
AVSiam Lin and Bertasius (2024) - ViT Radford et al. (2021) ViT Radford et al. (2021) $\times$
AVST Li et al. (2022) LSTM ResNet-18 He et al. (2016) VGGish Gemmeke et al. (2017)
CAT Ye et al. (2024) LLaMA2-7B Touvron et al. (2023) ImageBind Girdhar et al. (2023) ImageBind Girdhar et al. (2023) $\times$
ChatBridge Zhao et al. (2023) Vicuna-13B Chiang et al. (2023) ViT-G Sun et al. (2023) BEATs Chen et al. (2023a) $\times$
CIGN Mo et al. (2023) - ResNet-18 He et al. (2016) ResNet-18 He et al. (2016)
COCA Lao et al. (2023) Word Embedding ResNet-18 He et al. (2016) VGGish Gemmeke et al. (2017) $\times$
CONVLSTM Fayek and Johnson (2020) LSTM - Conv $\times$
CrossMAE Guo et al. (2024) - MAE He et al. (2022) AudioMAE Huang et al. (2022) $\times$
DCL Lv et al. (2023) DeBERTa-V3-Large He et al. (2021) ViT Radford et al. (2021) AST Gong et al. (2021)
DG-SCT Duan et al. (2023) - ViT Radford et al. (2021) HTS-AT Chen et al. (2022)
EEMC Wang et al. (2024a) RoBERTa Zhuang et al. (2021) ViT Radford et al. (2021) VGGish Gemmeke et al. (2017)
FCNLSTM Fayek and Johnson (2020) LSTM - Conv $\times$
GPT-4o Hurst et al. (2024) Transformer CLIP-ViT Transformer $\times$
GRU Antol et al. (2015) LSTM VGGNet Simonyan and Zisserman (2014) - $\times$
HCRN Le et al. (2020) BiLSTM ResNet-18 He et al. (2016) - $\times$
LAST-Att Liu et al. (2024b) LSTM Swin-Transformer-v2 Liu et al. (2022) Audio-Spectrogram-Transformer
LAVisH Lin et al. (2023) - ViT Radford et al. (2021) ViT Radford et al. (2021)
LAViT Yun et al. (2021) Transformer Vaswani et al. (2017) Transformer Vaswani et al. (2017) Transformer Vaswani et al. (2017)
LSTTA Liu et al. (2023) CLIP Radford et al. (2021) CLIP Radford et al. (2021) w2v-Conformer Gulati et al. (2020)
MAVEN Ma et al. (2025) Mixtral InternViT-300M-448px Chen et al. (2024b) Transformer $\times$
MCAN Yu et al. (2019) GloVe Pennington et al. (2014)+LSTM Faster R-CNN Ren et al. (2015) - $\times$
MCCD Ma et al. (2024) - - -
Meerkat Chowdhury et al. (2025) LLaMA2-7B Touvron et al. (2023) CLIP-ViT CLAP Elizalde et al. (2023)
OGM Wei et al. (2024) - ResNet-18 He et al. (2016) ResNet-18 He et al. (2016) $\times$
@@
OneLLM Han et al. (2024a) LLaMA2-7B Touvron et al. (2023) CLIP-ViT Unified Multimodal Encoder $\times$
OPM Wei et al. (2024) - ResNet-18 He et al. (2016) ResNet-18 He et al. (2016) $\times$
PSAC Li et al. (2019) Word Embedding CNN - $\times$
PSTP-Net Li et al. (2023) CLIP Radford et al. (2021) CLIP Radford et al. (2021) VGGish Gemmeke et al. (2017)
@@
QaP Liang et al. (2024) DeBERTa-V2-XLarge CLIP Radford et al. (2021) CLAP Elizalde et al. (2023) $\times$
Qwen2.5-VL Bai et al. (2025) MRoPE Bai et al. (2025) ViT Radford et al. (2021) - $\times$
RefAtomNet Peng et al. (2024) BERT ViT Radford et al. (2021) -
VALOR Liu et al. (2025a) BERT CLIP Radford et al. (2021) AST Gong et al. (2021) $\times$
VAST Chen et al. (2023b) BERT Devlin et al. (2019) ViT Dosovitskiy et al. (2020) BEATs Chen et al. (2023a) $\times$
VideoLLaMA-2 Cheng et al. (2024) Transformer CLIP Radford et al. (2021) BEATs Chen et al. (2023a)
VITA Fu et al. (2024) Mixtral Jiang et al. (2024) InternViT-300M-448px Chen et al. (2024b) CNN $\times$
Table 1: Architectural summary of representative Music AVQA methods. Each method lists the text, visual, and audio encoders used, along with an indication of whether explicit spatial-temporal (S-T) modeling is incorporated. Detailed descriptions of each method are provided in Appendix G and H.
@@ 4 Categorization of Music AVQA Methods Based on Architecture
CNN-based models. CNN-based methods utilize convolutional backbones such as ResNet or VGGish to encode modality-specific information into global or regional features, often relying on simpler late-stage fusion strategies. The AVST method exemplifies this approach, combining ResNet-18 visual embeddings and VGGish audio features through spatial attention modules to explicitly localize sound sources within visual frames. PSTP-Net extends this design by introducing a progressive refinement strategy that sequentially filters temporal segments and spatial regions, systematically narrowing down question-relevant audio-visual content prior to fusion. Although CNN-based models are computationally efficient and straightforward, their reliance on late fusion may pose challenges to capturing the complex temporal dynamics characteristic of musical performances.
Hybrid models. Hybrid models combine CNNs, transformers, and large language models (LLMs) to enable unified multimodal reasoning. They typically employ pre-trained encoders from both CNN and transformer families, integrated through sophisticated cross-modal fusion mechanisms. Representative examples include ChatBridge, CAT, OneLLM, and Meerkat. ChatBridge utilizes a perceiver-based multimodal transformer to merge modalities via language-aligned latent representations, followed by a frozen LLM for reasoning. CAT introduces modality-specific clue aggregation modules on top of ImageBind encodings, enabling precise question-driven multimodal grounding before passing information to a generative LLaMA2 LLM. OneLLM further generalizes multimodal integration by introducing a universal projection mechanism that allows a single LLM to interpret diverse modality embeddings seamlessly.
@@ 5 A Call on Specialized Multimodal Input Processing for Music AVQA
Compared to event-centric AVQA tasks that typically involve short, discrete sound events and lower-resolution recordings, Music AVQA deals with continuous, polyphonic streams spanning multiple spatial and temporal scales. Audio is commonly sampled at high rates (44.1 kHz or above) and often preserved in lossless formats to retain subtle timbral and articulatory detail Slizovskaia et al. (2021). Visual inputs tend to require higher resolution (1080p or above) and frame rates (30–60 fps) to capture nuanced performer motions such as bowing or fingering Gan et al. (2020); Jin et al. (2024). Even modest temporal offsets ($\approx$100–200ms) affect the perceived correspondence between gesture and sound. To improve synchronization and cue isolation, some recent models adopt preprocessing strategies like beat-based segmentation Böck et al. (2016) and harmonic-percussive separation Fitzgerald (2010), which can help surface rhythmically or acoustically meaningful content for downstream reasoning.
@@
Tokenization plays a central role in structuring inputs for multimodal reasoning, and recent Music AVQA models often tailor their strategies to preserve musical structure. For audio, models such as Amuse, DG-SCT, and PSTP-Net transform waveforms into Mel-spectrograms, which are then segmented via patch-based encoders like AST Gong et al. (2021) and HTS-AT or CNNs such as VGGish Gemmeke et al. (2017) and ResNet-18 He et al. (2016). Audio Flamingo Kong et al. (2024), for instance, uses overlapping 7-second windows in ClapCap Elizalde et al. (2024) to embed long-range audio context. Visual streams are frequently tokenized using ViT Radford et al. (2021) or Swin-based Liu et al. (2022) patch embeddings (e.g., in AVSiam Lin and Bertasius (2024) and LAVisH Lin et al. (2023)), while earlier models like AVST Li et al. (2022) use frame-level CNN features. Text tokenization is typically handled by subword models aligned with large language models (e.g., LLaMA2 Touvron et al. (2023), RoBERTa Zhuang et al. (2021)), as seen in ChatBridge Zhao et al. (2023) and OneLLM Han et al. (2024a). These tokenization schemes help preserve temporal granularity and modality alignment, which may be important for interpreting overlapping instruments, rhythmic changes, and localized visual cues.
@@
Music AVQA tasks often involve long-form performances with overlapping sources and evolving musical dynamics, which can create challenges for segmentation, annotation, and generalization. Unlike typical AVQA datasets centered on short clips and isolated actions, music-focused benchmarks (e.g., MUSIC-AVQA Li et al. (2022)) include multi-instrument performances spanning several minutes. These conditions place greater demands on dataset diversity to avoid overfitting to genre-specific patterns or ensemble configurations. To broaden coverage, some models are trained on data drawn from live performances, studio recordings, and synthetic renderings. However, the absence of symbolic structure can limit the model’s access to mid-level grounding. In this context, musically informed preprocessing (e.g., onset alignment, rhythmic segmentation, graph representation learning Huang et al. (2024a)) may support more interpretable and temporally aligned input representations.